OpenStudy (anonymous):

V =hS^2-hs^2, where h is the height of the prisms, S is the base length of the larger prism, and s is the base length of the smaller prism. the formula for the difference in volumes between two square prisms in factored form. h = 10cm, S = 30 cm, and s = 20 cm.

Directrix (directrix):

Volume Difference = h S² - hs² V.D. = h (S + s ) ( S - s)

Directrix (directrix):

h = 10cm, S = 30 cm, and s = 20 cm. V.D. = h (S + s ) ( S - s) V.D. = 10 ( 30 + 20)(30-20) V.D. = 5000 cubic centimers = 5 liters

OpenStudy (anonymous):

Oh, ok, I understand now. Thank you.

Directrix (directrix):

Glad to help.